OpenStudy (anonymous):

Determine the integral of x^2cosec x^3 cot x^3 dx

OpenStudy (dape):

Do the substitution u=x^3. Also the integral of csc(x)cot(x) is -csc(x).

OpenStudy (dape):

Proof of ∫csc(x)cot(x)=-csc(x)+C: \[ \int{\csc(x)\cot(x)\,dx}=\int{\frac{1}{\sin(x)}\frac{\cos(x)}{\sin(x)}\,dx} \] [Substitute u=sin(x), du=cos(x)dx] \[ = \int{\frac{du}{u^2}} =-\frac{1}{u}+C=-\frac{1}{\sin(x)}+C=-\csc(x)+C \]

OpenStudy (anonymous):

Ok great,thanx alot!!! :)